solve the equation by factoring.
2/3x^2+ 8 2/3x+20=0

Answer by MathTherapy(10549)

2/3x^2+ 8 2/3x+20=0
2x^2 + 26x + 60 = 0
(2x + 6)(x + 10) = 0
2x = - 6 or x + 10 = 0
x = - 3 or x = - 10
{- 3, - 10}